# RALAD: Bridging the Real-to-Sim Domain Gap in Autonomous Driving with Retrieval-Augmented Learning

### v2

@@ II-B Gap Between Real And Sim
In the field of autonomous driving, there exists a gap between simulation and reality [25], which is caused by discrepancies in lighting, textures, vehicle dynamics, and agent behaviors. To address this, researchers have developed two primary approaches: knowledge transfer learning and digital twins (DTs) [17, 28]. In knowledge transfer learning, the RG problem is compounded by uneven environmental sampling and complex physical parameters [26]. To overcome this, researchers have developed strategies such as curriculum learning, meta-learning, and domain randomization [27]. Domain randomization, in particular, helps align simulation parameters with real-world variability, facilitating the transfer of learned strategies to real-world applications. Conversely, digital twin technology creates virtual models of real-world entities or systems. A case in point is the SynFog dataset [28], which uses an end-to-end process to produce photo-realistic synthetic data. Nonetheless, despite their promise, these methods confront the issue of high computational expenses, particularly in complex, dynamic real-world settings.
@@ III-A Problem Formulation
The gap between real-world and simulation (real2sim) primarily arises from the cross-domain challenges between real and simulated data in autonomous driving. To address this, we consider two datasets: the Real dataset $D_{r}$ and the Sim dataset $D_{s}$. The Real dataset is defined as $D_{r}=\left\{X_{n},Y_{n}\right\}_{n=1}^{N_{r}}$, where $N_{r}$ is the number of images, and the Sim dataset is $D_{s}=\left\{\tilde{X_{m}},\tilde{Y_{m}}\right\}_{m=1}^{{N_{s}}}$, with ${N_{s}}$ being the number of simulated images, X and Y represent the image and ground truth, respectively. To bridge the gap between these two datasets, the first step is to establish a mapping relationship between them. We extract features from real and simulated images using an encoder function g(·), resulting in $H_{k}$ = g($X_{k}$) for real images and $\tilde{H}_{k}$ = g($\tilde{X}_{k}$) for simulated images. Our ultimate goal is to create pixel-to-pixel mapping relationship between features from different domain, thereby find the most similar real feature map from the real dataset for each simulated feature map. The formula is as follows:
@@ III-B Retrieval-Augmented Learning based on Optimal Transport in Real2Sim
To address the mapping relationship between features from different domains, we introduce global feature level Optimal Transport, because the previous RAL method was mainly local and focused on the differences between individual instances in high-definition medical images, the instance images were relatively clear and did not have much interference. However, in autonomous driving, the environment has a significant impact on vehicle recognition, such as in rainy or snowy weather. So we proposed a global feature level Optimal Transport to globally quantify the differences between reality and simulation. As shown in Figure 3, we use the encoder of the pre-trained model to extract features from a single simulated image and all real images, and then separate the extracted features from each image. Then the simulated feature values are subjected to OT calculations with all real feature values. This ensures that a pixel-to-pixel mapping relationship can be established for simulating features to find more perfect real features, as shown in Figure 5, where vehicles and lanes almost correspond. This mapping process is repeated on the simulated dataset until the traversal of the simulated dataset is completed.
@@ III-C Convex Merge and Fine-Tune
In the feature retrieval phase of autonomous driving, we have already obtained it through OT computed the optimal matching simulated feature $\tilde{H}^{*}$. Upon finding this match, we employ a convex merge operation to combine the real feature $H$ with the simulated feature $\tilde{H}^{*}$, resulting in a new composite feature $\hat{H}$, calculated as $\hat{H}=\pi(H,\tilde{H}^{*})$, as shown in Figure 2. The merging function $\pi$(·) is typically a convex combination method, this fusion also allows us to establish mapping relationships on features. The set of merged features denoted as $D_{m}=\{\hat{H_{i}}\}^{N_{m}}_{i=1}$, with $N_{m}$ being the number of features, is then utilized for the fine-tuning phase, aimed at enhancing the model’s ability to generalize from simulated to real data.
@@ IV-E Ablation Experiment
Convex Combination Ratio: In the feature merging stage of RALAD, different fusion ratios can produce varied results for the model. To explore this, we conducted comparison experiments with four different ratio settings, validating them on the KITTI and CARLA datasets, using a total of 1800 features. As shown in Table III, the 0.6:0.4 combination achieved a balanced performance, slightly decreasing on KITTI, but showing a substantial improvement on CARLA, making it an optimal trade-off between real and simulated data. In contrast, the 0.7:0.3 combination performed better on KITTI , but much worse on CARLA. As the CARLA ratio increased, performance on CARLA continued to improve, reaching the highest metrics with the 0.4:0.6 combination, but at the cost of reduced performance on KITTI, so we recommend and adopt this ratio for feature fusion to obtain better results.

### v3

@@ II-B Gap Between Real And Sim
In the field of autonomous driving, there exists a gap between simulation and reality [25], which is caused by discrepancies in lighting, textures, vehicle dynamics, and agent behaviors. To address this, researchers have developed two primary approaches: knowledge transfer learning and digital twins (DTs) [17, 28, 30]. In knowledge transfer learning, the reality gap problem is compounded by uneven environmental sampling and complex physical parameters [26]. To overcome this, researchers have developed strategies such as curriculum learning, meta-learning, and domain randomization [27]. Domain randomization, in particular, helps align simulation parameters with real-world variability, facilitating the transfer of learned strategies to real-world applications. Conversely, digital twin technology creates virtual models of real-world entities or systems. A case in point is the SynFog dataset [28], which uses an end-to-end process to produce photo-realistic synthetic data. Nonetheless, despite their promise, these methods confront the issue of high computational expenses, particularly in complex, dynamic real-world settings.
@@ III-A Problem Formulation
The gap between real-world and simulation (real2sim) primarily arises from the cross-domain challenges between real and simulated data in autonomous driving. To address this, we consider two datasets: the Real dataset $D_{r}$ and the Sim dataset $D_{s}$. The Real dataset is defined as $D_{r}=\left\{X_{n},Y_{n}\right\}_{n=1}^{N_{r}}$, where $N_{r}$ is the number of images, and the Sim dataset is $D_{s}=\left\{\tilde{X_{m}},\tilde{Y_{m}}\right\}_{m=1}^{{N_{s}}}$, with ${N_{s}}$ being the number of simulated images, X and Y represent the image and ground truth, respectively. To bridge the gap between these two datasets, the first step is to establish a mapping relationship between them. We extract features from real and simulated images using an encoder function g(·), resulting in $H_{k}$ = g($X_{k}$) for real images and $\tilde{H}_{k}$ = g($\tilde{X}_{k}$) for simulated images. Our ultimate goal is to create mapping relationship between features from different domain, thereby find the most similar real feature map from the real dataset for each simulated feature map. The formula is as follows:
@@ III-B Retrieval-Augmented Learning based on Optimal Transport in Real2Sim
To address the mapping relationship between features from different domains, we introduce global feature level Optimal Transport, because the previous RAL method was mainly local and focused on the differences between individual instances in high-definition medical images, the instance images were relatively clear and did not have much interference. However, in autonomous driving, the environment has a significant impact on vehicle recognition, such as in rainy or snowy weather. So we proposed a global feature level Optimal Transport to globally quantify the differences between reality and simulation. As shown in Figure 3, we use the encoder of the pre-trained model to extract features from a single simulated image and all real images, and then flatten the feature map to generate row vectors. Then the simulated vectors are subjected to OT calculations with all real vectors. This ensures that a distribution correlation can be established between simulated and real features, enabling the discovery of more optimal real features, as shown in Figure 5, where vehicles and lanes almost correspond. This mapping process is repeated on the simulated dataset until the traversal of the simulated dataset is completed.
@@ III-C Convex Merge and Fine-Tune
In the feature retrieval phase of autonomous driving, we have already obtained it through OT computed the optimal matching simulated feature $\tilde{H}^{*}$. Upon finding this match, we employ a convex merge operation to combine the real feature $H$ with the simulated feature $\tilde{H}^{*}$, resulting in a new composite feature $\hat{H}$, calculated as $\hat{H}=\pi(H,\tilde{H}^{*})$, as shown in Figure 2. The merging function $\pi$(·) is typically a convex combination method, convex fusion of features H and ${H}^{*}$ is achieved through complementary weight coefficients, where the coefficients ensure that the fusion result is located on the line segment formed by the two features, achieving linear interpolation and balance in the feature space, this fusion also allows us to establish mapping relationships on features. The set of merged features denoted as $D_{m}=\{\hat{H_{i}}\}^{N_{m}}_{i=1}$, with $N_{m}$ being the number of features, is then utilized for the fine-tuning phase.
@@ IV-E Ablation Experiment
Convex Combination Ratio: In the feature merging stage of RALAD, different fusion ratios can produce varied results for the model. To explore this, we conducted comparison experiments with four different ratio settings, validating them on the KITTI and CARLA datasets, using a total of 1800 features. As shown in Table III, KITTI: CARLA represent the weight coefficients for real and simulated feature maps, respectively, during the merge, the ratio of the convex merge is based on the  [18], and we have also conducted additional validation experiments, which demonstrate that the 0.6:0.4 combination achieved a balanced performance, slightly decreasing on KITTI, but showing a substantial improvement on CARLA, making it an optimal trade-off between real and simulated data. In contrast, the 0.7:0.3 combination performed better on KITTI , but much worse on CARLA. As the CARLA ratio increased, performance on CARLA continued to improve, reaching the highest metrics with the 0.4:0.6 combination, but at the cost of reduced performance on KITTI, so we recommend and adopt this ratio for feature fusion to obtain better results.
@@ V CONCLUSIONS
VI ACKNOWLEDGEMENT
This work was supported in part by the National Natural Science Foundation of China (62072321), the Science and Technology Program of Jiangsu Province (BZ2024062), the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (22KJA520007), Suzhou Planning Project of Science and Technology (2023ss03).